OpenStudy (anonymous):

What is the equation of the line in slope-intercept form that passes through the point (–2, 4) and is perpendicular to the line y = –2x + 4? a. y = -1/2x – 5 b. y=-1/2x-2 c. y=1/2x+2 d. y=1/2x+5

OpenStudy (anonymous):

Whats slope of y = –2x + 4?

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

So can you find the perpendicular slope?

OpenStudy (anonymous):

wouldnt that be 1/2

OpenStudy (anonymous):

Correct :) Now after you have the perpendicular slope, plug m = 1/2 and (–2, 4) into point slope form formula:

OpenStudy (anonymous):

so would it be y=1/2x+2?

OpenStudy (anonymous):

well im confused.

OpenStudy (anonymous):

yeah, i figured it out and got y=1/2x+5

OpenStudy (anonymous):

Correct answer! You won't be confused if you know the steps to proceed !

OpenStudy (anonymous):

thanks!